# LLMTaxo: Leveraging Large Language Models for Constructing Taxonomy of Factual Claims from Social Media

### v1

@@ Abstract
With the vast expansion of content on social media platforms, analyzing and comprehending online discourse has become increasingly complex. This paper introduces $\mathsf{LLMTaxo}$, a novel framework leveraging large language models for the automated construction of taxonomy of factual claims from social media by generating topics from multi-level granularities. This approach aids stakeholders in more effectively navigating the social media landscapes. We implement this framework with different models across three distinct datasets and introduce specially designed taxonomy evaluation metrics for a comprehensive assessment. With the evaluations from both human evaluators and GPT-4, the results indicate that $\mathsf{LLMTaxo}$ effectively categorizes factual claims from social media, and reveals that certain models perform better on specific datasets.
@@ 1 Introduction
Misinformation, also known as false or misleading information Wu et al. (2019), has the potential to sway public perception, cause confusion, and influence people’s decision-making processes Del Vicario et al. (2016); Muhammed T and Mathew (2022). Social media platforms, in particular, facilitate the rapid sharing of vast amounts of content, blending accurate information with falsehoods Allcott et al. (2019). Social media’s global reach and ease of use have transformed how millions of users exchange opinions, news, and factual claims in real-time, making it fertile ground for misinformation Aïmeur et al. (2023). Factual claims, which are assertions that can be verified as either true or false, are a common vehicle for misinformation Ni et al. (2024). These claims, whether accurate or not, have a profound societal impact, as the public tends to believe a factual claim is true regardless of its truthfulness Moravec et al. (2018); Ognyanova et al. (2020); Xiao et al. (2021); Zhu et al. (2022); Zhang et al. (2024a).
@@
The dynamic nature of social media often leads to repetitive or reformulated claims, complicating the identification and validation of factual content Zhou et al. (2015), making it difficult for non-technical individuals and researchers from different fields to navigate the vast amounts of data Suarez-Lledo and Alvarez-Galvez (2021); Hook and Verdeja (2022); Muhammed T and Mathew (2022). Therefore, there is an urgent need for scalable, automated tools to organize and analyze factual claims in a systematic way, enabling stakeholders such as researchers, fact-checkers, and policymakers to better navigate this information landscape. For example, Tambini (2017) categorized fake news into different types to describe a range of fake news phenomena. Kumar and Shah (2018) constructed taxonomies for false information on social media platforms based on different characteristics, such as with the intention to deceive or not and opinion-based or fact-based. Zhao and Tsang (2022) proposed a taxonomy of misinformation on social media based on falsity level and evidence type. In this paper, we present a novel framework, $\mathsf{LLMTaxo}$, for the automated construction of a taxonomy of factual claims on social media by using large language models (LLMs) to generate multi-level topics for each claim.
@@
A taxonomy for organizing information assets starts with broad categories and branches into increasingly specific subcategories. In this way, factual claims can be categorized into meaningful categories, allowing for the identification of distinct claims, reduction of redundancy, and exploration of information at multiple levels of granularity. For example, broad categories can group COVID-19 vaccine-related claims into overarching topics such as public health, while more detailed categories can address specific claims about vaccine safety. Such structured organization enhances fact-checking and research workflows, and offers a clearer understanding of the themes and patterns in factual claims.
Our framework, $\mathsf{LLMTaxo}$, clusters semantically similar factual claims, identifies distinct ones, and generates topics at multi-granularities, thus organizing claims into a hierarchical taxonomy with broad, medium, and detailed topics. The LLMs have rich background knowledge that is not only applicable now but also in the future while the online content changes over time. By leveraging the advanced capabilities of LLMs and few-shot learning, the framework minimizes human involvement and automates the time-intensive task of categorizing claims while maintaining adaptability to the evolving nature of social media discourse.
We addressed multiple challenges presented in developing $\mathsf{LLMTaxo}$. One key challenge lies in the semantic variability of claims, where similar ideas are conveyed differently. $\mathsf{LLMTaxo}$ addresses this by clustering and identifying distinct claims. Another challenge is the scalability and applicability of the framework across diverse datasets and domains. To tackle this, we incorporate few-shot in-context learning for LLMs, enabling them to be easily adapted to large datasets and various domains. To demonstrate the generalizability of $\mathsf{LLMTaxo}$, we evaluated it using carefully designed metrics across three distinct datasets from different topics and data sources, including posts from X (formerly Twitter) and Facebook, covering topics such as COVID-19 vaccines, climate change, and cybersecurity.
The evaluation results demonstrate the effectiveness of $\mathsf{LLMTaxo}$ in enhancing our understanding of the social media landscape. By identifying distinct factual claims, the framework significantly reduces redundancy. The hierarchical taxonomy produced by $\mathsf{LLMTaxo}$ enables a multi-level analysis of claims, allowing for exploration at varying levels of detail. Furthermore, the framework exhibits strong performance across diverse datasets, showcasing its adaptability and potential for broad application. The datasets and codebase are available at https://anonymous.4open.science/r/LLMTaxo-2595.
This paper presents several key contributions:
@@
  • We introduce the first taxonomy of factual claims on social media constructed using LLMs. This resource can be directly integrated into fact-checking workflow and various research fields.
  • We are the first to utilize LLMs for generating topics at multiple granularities.
@@
  • We develop a set of evaluation metrics for taxonomy and claim-topic pairs to comprehensively assess the quality of taxonomy.
@@ 2 Related Work
Although taxonomy construction has been extensively studied, the definitions of specific problems vary. Generally, taxonomies are hierarchically structured classifications of concepts, terms, and entities that help users organize, retrieve, and navigate information Carrion et al. (2019); Yang (2012). Generic taxonomy construction tasks typically involve short concept terms or entity names, often represented as hypernym-hyponym pairs Zhang et al. (2018); Huang et al. (2020).
@@
The rise of LLMs has significantly advanced many natural language processing tasks, but limited effort has been devoted to taxonomy construction. Chen et al. (2020) employed pretrained language models to construct taxonomic trees, while Chen et al. (2023) compared prompting and fine-tuning approaches for hypernym taxonomy construction. Additionally, Shah et al. (2023); Wan et al. (2024) introduced end-to-end pipelines that integrate LLMs to generate, refine, and apply labels for user intent analysis in log data.
@@
LLMs have emerged as a promising alternative to traditional topic generation approaches. They can generate topics from a given set of documents without requiring predefined labels or training data Sarkar et al. (2023). This capability allows for more flexible and context-aware topic extraction. For instance, Mu et al. (2024) introduced a framework that prompts LLMs to generate topics and adhere to human guidelines for refining and merging topics. Recent research has also explored different LLM-based topic modeling techniques. For example, BERTopic Grootendorst (2022) has shown superior performance in terms of diversity and coherence across multiple datasets Jung et al. (2024). Yet, to the best of our knowledge, no existing work has generated topics from multi-granularities.
@@ 3 Methodology
Analyzing social media data presents numerous challenges, including the overwhelming volume of content with high repetition and the substantial human effort required to explore the data. Our $\mathsf{LLMTaxo}$ framework is designed to systematize the chaotic nature of social media through the automated construction of a taxonomy of factual claims. This taxonomy serves as a hierarchical classification system, enhancing the accessibility and navigability of information for users by categorizing claims into broad, medium, and detailed topics. The $\mathsf{LLMTaxo}$ framework initially identifies factual claims from social media posts, subsequently clusters similar claims to discern and select distinct ones, thereby reducing redundancy. It then leverages LLMs to generate topics for each distinct claim at multiple levels of granularity, ultimately constructing a structured taxonomy. This structured taxonomy provides an organized way to manage and explore factual claims from social media platforms. The overview of our framework is shown in Figure 1. In this section, we first explain how we utilize LLMs to develop a taxonomy. We then describe the techniques employed for detecting factual claims and the methods used to identify and select distinct claims.
@@ 3.1 Taxonomy Construction
The primary goal of this study is to construct a hierarchical taxonomy. The hierarchical design is informed by the prototype theory Geeraerts (2006), which guides cognitive categorization. Formally, a hierarchical taxonomy $\mathcal{T}=(T_{b},T_{m},T_{d},f_{m},f_{d})$, that organizes a collection of factual claims $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$ into topics at multiple levels of granularity, ranging from broader to more fine-grained. Here, $f_{m}$ and $f_{d}$ define the hierarchical relationships between topics, where $f_{m}:T_{m}\rightarrow T_{b}$ maps each medium topic to its corresponding broad topic, and $f_{d}:T_{d}\rightarrow T_{m}$ maps each detailed topic to its respective medium topic. Specifically, the taxonomy was constructed using a three-tiered structure consisting of broad topic $t_{b}\in T_{b}$ (representing the general theme), medium topic $t_{m}\in T_{m}$ (reflecting intermediate distinctions), and detailed topic $t_{d}\in T_{d}$ (highlighting finer aspects) for each claim. Each claim $c\in C$ is assigned topics at three levels: $\phi(c)=(t_{b},t_{m},t_{d})$, where $\phi$ is a function that maps each claim $c$ to a tuple containing its broad, medium, and detailed topics. For example, consider a factual claim $c$ with $\phi(c)=(t_{b},t_{m},t_{d})$ related to COVID-19 vaccine: “Myocarditis is up TEN times due to the Covid Vaccine…” The broad topic $t_{b}$ is Vaccine Safety and Effectiveness, the medium topic $t_{m}$ is Vaccine Side Effects, and the detailed topic $t_{d}$ is Myocarditis Side Effect, as shown on the right side of Figure 1. The hierarchical structure ensures that the taxonomy captures both high-level themes and more nuanced distinctions across factual claims. To automate the taxonomy construction process, we create learning examples that contain both factual claims and their corresponding topics. The topics from the learning examples form a seed taxonomy, which serves as a foundation for LLM-based expansion. We prompt the LLMs with both the learning examples and the seed taxonomy to generate topics for each distinct factual claim. After generating topics for all claims, we consolidate them to construct a refined taxonomy, as shown in Figure 1.
@@ 3.1.1 Learning Examples and Seed Taxonomy
While LLMs possess broad general knowledge, they are not inherently equipped to generate hierarchical topics for factual claims across different granularities specific to our needs. Our initial experiments indicated that LLMs often produce inconsistent topics for similar claims, leading to an unwieldy number of categories that hinder comprehension, as further detailed in Section 4.4. To mitigate this issue, we propose to incorporate a seed taxonomy $\mathcal{S}$ derived from a set of sample factual claims. This initial taxonomy aids the LLMs by providing a foundation to expand upon. We manually create $k$ learning examples =$\mathcal{L}=\{(c_{1},t_{b_{1}},t_{m_{1}},t_{d_{1}}),...,(c_{k},t_{{b}_{k}},t_{m_{k}},t_{{d}_{k}})\}$, representing claims with their respective topics at each level of granularity. The topics from the $k$ claims form the seed taxonomy, denoted as $\mathcal{S}=\{(t_{b_{1}},t_{m_{1}},t_{d_{1}}),...,(t_{b_{k}},t_{m_{k}},t_{d_{k}})\}$. These learning examples and the seed taxonomy facilitate few-shot in-context learning, helping to stabilize the variation and number of topics generated. The size $k$ of learning examples varies across datasets.
To compile the learning examples and the seed taxonomy, we randomly select a subset $\mathcal{R}\subset\mathcal{C}$ containing $m$ distinct claims for annotation. Recognizing that direct human annotation is labor-intensive and the annotation requires high accuracy and consistency, we employ GPT-3.5 Brown et al. (2020) to assist in topic annotation for these claims.
Each claim $c_{i}\in\mathcal{R}$ is processed by the LLMs, which is prompted to generate topics at three levels of granularity. After evaluating the LLM-generated topics, we refine and annotate each claim $c_{i}$ with $(t_{b_{i}},t_{m_{i}},t_{d_{i}})$. The LLM-guided process reduces manual effort and allows for tailored annotations based on specific requirements, which is adaptable across various domains. Below is an example of the prompt used for the COVID-19 vaccine dataset:
@@
To achieve consistency and minimize topic variation, we restrict the number of topics and assign identical topics to claims within the same category. Note that no all factual claims have three-level topics, as some may be too brief or ambiguous to be categorized distinctly into broad, medium, and detailed topics. In such cases, the claim may align with only one or two topic levels. Upon annotating all $m$ claims, we reviewed the most frequently occurring broad topics and established representative sets. The topics from the final set of $k$ factual claims form the seed taxonomy $\mathcal{S}$. The $k$ claim-topic pairs are used as learning examples $\mathcal{L}$. The seed taxonomy and the learning examples guide the LLMs to generalize the task of topic generation across a diverse range of claims while ensuring consistent and structured outputs.

### v2

@@ Abstract
With the rapid expansion of content on social media platforms, analyzing and comprehending online discourse has become increasingly complex. This paper introduces $\mathsf{LLMTaxo}$, a novel framework leveraging large language models for the automated construction of taxonomies of factual claims from social media by generating topics at multiple levels of granularity. The resulting hierarchical structure significantly reduces redundancy and improves information accessibility. We also propose dedicated taxonomy evaluation metrics to enable comprehensive assessment. Evaluations conducted on three diverse datasets demonstrate $\mathsf{LLMTaxo}$’s effectiveness in producing clear, coherent, and comprehensive taxonomies. Among the evaluated models, GPT-4o mini consistently outperforms others across most metrics. The framework’s flexibility and low reliance on manual intervention underscore its potential for broad applicability.
@@ 1 Introduction
Misinformation, also known as false or misleading information Wu et al. (2019), has the potential to sway public perception, cause confusion, and influence people’s decision-making processes Del Vicario et al. (2016); Muhammed T and Mathew (2022). Social media platforms, in particular, facilitate the rapid sharing of vast amounts of content, blending accurate information with falsehoods Allcott et al. (2019). Social media’s global reach and ease of use have transformed how millions of users exchange opinions, news, and factual claims in real-time, making it fertile ground for misinformation Aïmeur et al. (2023). Factual claims, which are assertions that can be verified as either true or false, are a common vehicle for misinformation Ni et al. (2024). These claims, whether accurate or not, have a profound societal impact, as the public tends to believe a factual claim is true regardless of its truthfulness Moravec et al. (2018); Ognyanova et al. (2020); Xiao et al. (2021); Zhu et al. (2022); Zhang et al. (2024a); Zhu et al. (2025).
@@
The dynamic nature of social media often leads to repetitive or reformulated claims, complicating the identification and validation of factual content Zhou et al. (2015), making it difficult for non-technical individuals and researchers from different fields to navigate the vast amounts of data Suarez-Lledo and Alvarez-Galvez (2021); Hook and Verdeja (2022); Muhammed T and Mathew (2022). This challenge calls for scalable and automated tools to systematically organize and analyze factual claims, helping stakeholders—such as researchers and fact-checkers—more effectively navigate the complex information landscape.
One promising approach to addressing this challenge is the use of taxonomies, which provide structured systems for organizing complex information. Taxonomy has been applied in misinformation categorization. For example, Tambini (2017) classified fake news into distinct categories to capture the diverse forms and manifestations of the phenomenon. Kumar and Shah (2018) constructed taxonomies of false information on social media based on various characteristics, such as whether the content is intended to deceive and whether it is opinion-based or fact-based. Zhao and Tsang (2022) proposed a taxonomy of misinformation on social media based on falsity level and evidence type. Yet, these approaches do not construct topic taxonomies, which are particularly valuable for exploring and organizing factual claims. In practice, many fact-checking websites, such as PolitiFact, [^1] https://www.politifact.com/ Snopes, [^2] https://www.snopes.com/ and Full Fact, [^3] https://fullfact.org/ use topics to organize fact checks. For example, the claim “Right to Try experimental drug program saved thousands and thousands of lives” in PolitiFact is given topics such as “Elections” and “Health Care.” However, these topics are too broad to help users efficiently locate specific claims.
@@
Motivated by the challenges, prior studies, and practical applications discussed above, this paper introduces $\mathsf{LLMTaxo}$, a novel framework utilizing large language models (LLMs) for automatically constructing a topic taxonomy given a collection of factual claims from social media within a specific topic domain. The taxonomy starts with broad categories and branches into increasingly specific subcategories. In this way, factual claims can be categorized into hierarchical categories, allowing for the efficient exploration of information at multiple levels of granularity. For example, broad categories can group COVID-19 vaccine-related claims into overarching topics such as “Public Health,” while more detailed categories can address specific claims about “Vaccine Safety and Effectiveness.”
Our framework, $\mathsf{LLMTaxo}$, organizes factual claims into a coherent hierarchical taxonomy by clustering semantically similar claims, identifying distinct ones, and generating topics at multiple levels of granularity—broad, medium, and detailed. Leveraging the rich background knowledge of LLMs and few-shot learning, $\mathsf{LLMTaxo}$ minimizes human involvement and automates the categorization process. It effectively addresses challenges such as semantic variability (the expression of similar ideas in diverse ways) and ensures scalability and adaptability to social media discourse across various datasets and topic domains.
To demonstrate the generalizability of $\mathsf{LLMTaxo}$, we conducted evaluations using carefully designed metrics across three distinct datasets spanning different topic domains and data sources. These include social media posts from X (formerly Twitter) and Facebook, covering domains such as COVID-19 vaccines, climate change, and cybersecurity. The evaluation results, showing all models scoring above 3.0 out of 5 across all metrics, with a highest score of 4.9, demonstrate the effectiveness of $\mathsf{LLMTaxo}$ in generating clear and comprehensive taxonomies. By identifying distinct factual claims, the framework substantially reduces redundancy. The hierarchical structure of the taxonomy produced by $\mathsf{LLMTaxo}$ allows users to explore claims at multiple levels of detail. Moreover, $\mathsf{LLMTaxo}$ exhibits consistently strong performance across diverse datasets, demonstrating its adaptability and potential for wide-ranging applications.
In summary, this paper makes several key contributions:
@@
  • We introduce the first multi-level taxonomy of factual claims from social media constructed using LLMs. This taxonomy can be integrated into fact-checking workflows and applied across various research domains.
  • We are the first to utilize LLMs for generating multi-granularity topics.
@@
  • We develop a set of evaluation metrics for comprehensively assessing the structural quality of taxonomies and their claim-topic alignment.
@@
  • We release a curated dataset and a public codebase at https://github.com/idirlab/LLMTaxo. They are valuable for supporting both future research and reproducibility.
@@ 2 Related Work
Although taxonomy construction has been extensively studied, the definitions of specific problems vary. Generally, taxonomies are hierarchically structured classifications of concepts, terms, and entities that help users organize, retrieve, and navigate information Carrion et al. (2019); Yang (2012). Generic taxonomy construction tasks typically involve short concept terms or entity names, which are often structured as hypernym-hyponym pairs Zhang et al. (2018); Huang et al. (2020).
@@
The rise of LLMs has significantly advanced many natural language processing tasks, but limited effort has been devoted to taxonomy construction. Chen et al. (2020) employed pretrained language models to construct taxonomic trees, while Chen et al. (2023) compared prompting and fine-tuning approaches for hypernym taxonomy construction. Additionally, Shah et al. (2023) and Wan et al. (2024) introduced end-to-end pipelines that integrate LLMs to generate, refine, and apply labels for user intent analysis in log data. Our work extends this line of research by applying LLMs to generate multi-level taxonomies for factual claims and social media posts. Like prior efforts, we harness LLMs’ generative and semantic capabilities, but we focus on taxonomy construction in open-domain, noisy, and high-variance contexts where entity boundaries and hierarchical relationships are less clear.
@@
LLMs have emerged as a promising alternative to traditional topic generation approaches. They can generate topics from a given set of documents without requiring predefined labels or training data Sarkar et al. (2023). This capability allows for more flexible and context-aware topic extraction. For instance, Mu et al. (2024) introduced a framework that prompts LLMs to generate topics and adhere to human guidelines for refining and merging topics. Recent research has also explored different LLM-based topic modeling techniques. For example, BERTopic Grootendorst (2022) has shown superior performance in terms of diversity and coherence across multiple datasets Jung et al. (2024). Yet, to the best of our knowledge, no existing work has generated topics at multiple levels of granularity.
@@ 3 Methodology
Analyzing social media data presents numerous challenges, including the overwhelming volume of content with high repetition and the substantial human effort required to explore the data. Our $\mathsf{LLMTaxo}$ framework is designed to systematize online content through automated construction of a taxonomy of factual claims. This taxonomy serves as a hierarchical classification system. It enhances the accessibility and navigability of information for users by categorizing claims into broad, medium, and detailed topics.
$\mathsf{LLMTaxo}$ initially identifies factual claims from social media posts, and subsequently clusters similar claims to discern and select distinct ones, thereby reducing redundancy. It then leverages LLMs to generate topics for each distinct claim at multiple levels of granularity, ultimately constructing a structured taxonomy. The overview of our framework is shown in Figure 1.
@@ 3.1 Taxonomy Construction
[…]
Note that the same topic label may appear under different parent topics, but these instances represent conceptually distinct topics. For example, the medium topic “Vaccine Safety” under broad topic “Public Opinion” is essentially “Public Opinion about Vaccine Safety,” which is meaningfully different from “Vaccine Safety” under “Government Policies.” To minimize redundancy while preserving clarity, we use concise topic labels for medium and detailed topics, as their inherent meanings are informed by their parent topics.
Given a collection of factual claims $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$, each claim $c\in C$ is assigned a tuple of topics at three levels—broad, medium, and detailed—with a mapping function $\phi$. More formally, $\phi(c)=(t_{b},t_{m},t_{d})$ where $t_{b}\in T_{b}$, $t_{m}\in T_{m}$, and $t_{d}\in T_{d}$. Note that the detailed topic is optional because some claims may be too brief or ambiguous to be categorized distinctly at all three levels.
For example, the right side of Figure 1 illustrates the topic tuple for a claim related to COVID-19 vaccine: “Myocarditis is up TEN times due to the Covid Vaccine…” The broad topic $t_{b}$ is “Vaccine Safety and Effectiveness,” the medium topic $t_{m}$ is “Vaccine Side Effects,” and the detailed topic $t_{d}$ is “Myocarditis Side Effect.”
To automate the process of taxonomy construction given a collection of factual claims $\mathcal{C}$, we create learning examples that contain both sample claims and their corresponding topics. The topics from the learning examples form a seed taxonomy, which serves as a foundation for LLM-based expansion. To generate topics for each distinct factual claim in $\mathcal{C}$, we prompt the LLMs with both the learning examples and the seed taxonomy. After generating topics for all claims, we consolidate them to construct a refined taxonomy. This process is illustrated in Figure 1.
@@ 3.1.1 Learning Examples and Seed Taxonomy
While LLMs possess broad general knowledge, they are not pretrained to directly generate hierarchical topics for factual claims at different granularities. Our initial experiments showed that LLMs often generate inconsistent topics for similar claims, resulting in an excessive and unwieldy number of topics that hinder comprehension, as further detailed in Section 4.4. To mitigate this issue, we propose to expand and refine a seed taxonomy $\mathcal{S}$ derived from a set of sample factual claims. This initial taxonomy aids the LLMs by providing a foundation to expand upon. We leverage an LLM with human-in-the-loop supervision to create $k$ learning examples $\mathcal{L}$$\{\langle c_{1},(t_{b}^{(1)},t_{m}^{(1)},t_{d}^{(1)})\rangle,...,\langle c_{k},(t_{b}^{(k)},t_{m}^{(k)},t_{d}^{(k)})\rangle\}$, representing claims with their respective topics at each level of granularity. The 3-level topic tuples from the $k$ claims form the seed taxonomy, denoted as $\mathcal{S}$$\{(t_{b}^{(1)},t_{m}^{(1)},t_{d}^{(1)}),...,(t_{b}^{(k)},t_{m}^{(k)},t_{d}^{(k)})\}$. These learning examples and the seed taxonomy facilitate few-shot in-context learning, helping to stabilize the variation and number of topics generated. The size $k$ of learning examples varies across datasets.
To compile the learning examples and the seed taxonomy, we randomly select a subset $\mathcal{R}\subset$$\mathcal{C}$ containing $m$ distinct claims for annotation. Recognizing that direct human annotation is labor-intensive and the annotation requires high accuracy and consistency, we employ GPT-3.5 Brown et al. (2020) to assist human in topic annotation for these claims.
Each claim $c\in\mathcal{R}$ is processed by an LLM prompted to generate topics at three levels of granularity. Below is an example of the prompt used for the COVID-19 vaccine topic domain.
@@
Examples: ...
$\mathsf{LLMTaxo}$ adopts a multi-round human-in-the-loop annotation process. In each round, the LLM generates topics for a batch of 10 claims. These topics are then evaluated and refined by annotators, resulting in a finalized annotation of each claim $c$ with topic tuple $(t_{b},t_{m},t_{d})$, forming a claim-topic pair $\langle c,(t_{b},t_{m},t_{d})\rangle$. The LLM may produce diverse topic labels in the initial run. To ensure consistency and minimize topic variation, we restrict the number of distinct topics and standardize topic labels across similar claims during the refinement process. For example, if the LLM generates “COVID-19 Vaccine Mandates” for claim A and “Vaccine Mandates” for claim B, the annotators are expected to retain a single, generalized topic (e.g., “Vaccine Mandates”) for both claims.
The refined claim-topic pairs are subsequently used as in-context examples to guide the LLM in the next round. This process is iterated until topics are generated for all $c\in\mathcal{R}$. The human-in-the-loop process reduces manual effort and allows for tailored annotations based on specific requirements, which is adaptable across various topic domains.
Upon annotating all $m$ claims, the annotators review the most frequent broad topics and manually pick $k$ representative claim-topic pairs as the learning examples $\mathcal{L}$. The topics from these $k$ examples form the seed taxonomy $\mathcal{S}$. The seed taxonomy and the learning examples guide the LLMs to generalize the task of topic generation across a diverse range of claims while ensuring consistent and structured outputs.